A sample of nitrogen gas is condensed in a piston with a feely moving cylinder, allowing the volume to vary. At 0.0C, the volume of the gas is 375mL. To what temperature must the gas be heated to occupy a volume of 500.0mL


1
Expert's answer
2021-09-26T12:06:43-0400

"\frac{V_1}{T_1}=
\frac{V_2}{T_2}"


"\frac{375}{273}=
\frac{500}{T_2}"


"T_2=\frac{273×500}{375}=364K =91°C"